# LlaSMol: Advancing Large Language Models for Chemistry with a Large-Scale, Comprehensive, High-Quality Instruction Tuning Dataset

### v1

@@ Appendix A Details of Experimental Setup
In this section, we introduce the details of our experimental setups, including the training and inference details of our LlaSMol models and the compared models. We also give detailed explanations of the metrics used in Section 5.3, as well extra metrics that we will use in Appendix B.
A.1 LlaSMol Models
@@
The base models used for developing LlaSMol are Galactica444https://huggingface.co/facebook/galactica-6.7b (Taylor et al., 2022), Llama 2555meta-llama/Llama-2-7b-hf (Touvron et al., 2023b), Code Llama666codellama/CodeLlama-7b-hf (Roziere et al., 2023) and Mistral777mistralai/Mistral-7B-v0.1 (Jiang et al., 2023). We conduct instruction tuning on our SMolInstruct, and the resulting models are called named as LlaSMol, LlaSMol, LlaSMol, and LlaSMol, respectively. Expect for being based on different base models, their training and evaluation configurations are identical, as described as follows.
@@
During inference, we adopt beam search as the generation strategy for simplicity. Due to the need of evaluations on the top- predicted answers (as in Appendix B, where varies for different tasks, we generate different numbers of sequences for different tasks by setting the num_return_sequences argument in the Huggingface Transformers library (Wolf et al., 2020). Specifically, it is set to 5 for NC-I2S, NC-S2I, FS, and MG; 3 for NC-I2F and NC-S2F; 1 for all the PP tasks; and 10 for RS. The beam size is set to num_return_sequences + 3 for all the tasks. The maximum number of new generated tokens is set to 1024.
A.2 Compared Models
We introduce each of the compared models in details, including their training (if applicable) and inference process.
A.2.1 GPT-4
@@
We conduct both -shot evaluations, where is the number of provided ICL examples. For -shot evaluation, the ICL part in the template is removed from the queries. In -shot evaluation, for each sample,the ICL examples are randomly selected from the training set. The results of these settings are shown in Appendix B, which reals that these settings’ performance is not consistent across all the tasks. Since -shot shows the best performance on most tasks, we report its results in Section 5.3.
In the evaluations, we use the default generation strategy set in the API. To generate the same number of results for each sample (as described in Section A.1), we set the argument n in the API, which controls the number of output sequences.
@@
A.2.2 Galactica
@@
Galactica (Taylor et al., 2022) is a LLM without undergoing instruction tuning. To evaluate it on SMolInstruct, we follow the instructions in the paper (Taylor et al., 2022) and the repository888https://github.com/paperswithcode/galai to create the queries for each task. We use zero-shot setting, as its official instruction does not suggest using few-shot setting. The generation configuration is set identical to that of our LlaSMol models (Section A.1).
@@
Galatica’s outputs may contain extra text other than the expected answers. Therefore, with heuristic rules and regular expression matching, we implement a program to extract the answers from the outputs of the models.
A.2.3 Llama 2, Code Llama, and Mistral
@@
For our base models (Llama 2, Code Llama, and mistral), since they are not trained on SMolInstruct and have not seen the diverse queries in the dataset, we use the same query templates as those used for GPT-4 (Section A.2.1). We use the one-shot setting for them, as it would improve models’ abiltity to follow the instructions and generate answers in a more formated way. In addition, the generation configuration (including beam size, output sequence numbers, etc) is set identical to that of our LlaSMol models (Section A.1).
@@
Although we try our best to make the output format as clear as possible in the queries, these three models still cannot follow the instructions and their outputs are in various formats. By heuristic rules and regular expression matching, we implement a program to extract the answers from the outputs of each of the models.
A.2.4 Molinst
@@
Molinst is a Llama 2 model fine-tuned on Mol-Instructions (Fang et al., 2023). On the shared tasks between Mol-Instructions and SMolInstruct (including MC, MG, FS, and RS), we directly use the query templates from Mol-Instructions to achieve better results. On other tasks, we create one query template for each task following the style of Mol-Instructions. We use zero-shot on its evaluation, as Mol-Instructions does not contain any one-shot use case.
@@
The outputs of Molinst may also contain extra text other than the expected answers, especially on its unseen tasks. Thus, we also implement a program to extract the answers.
A.3 Task-Specific, Non-LLM Based SoTA Models
A.3.1 STOUT for NC-I2S and NC-S2I
@@
STOUT is a encoder-decoder model trained on SMILES and IUPAC name paired data, and it is capable of conducting the NC-2iS and NC-S2I tasks. Due to the lack of training code, we cannot re-train it on our dataset, and directly use their released model checkpoint999https://github.com/Kohulan/Smiles-TO-iUpac-Translator. Since it may have encounter some test samples of SMolInstruct during training, the evaluation results in Table 2 may be higher than its real performance.
A.3.2 RDKit for NC-S2F
@@ A.4 Evaluation Metrics
Fingerprint-based Tanimoto Similarity (FTS). It is an important metric type commonly used in cheminformatics. It measures the structural similarity between molecules. The one we report in Section 5.3 is one of this type, called Morgan FTS, which leverages Morgan method to calculate the fingerprint (Morgan, 1965).
@@
Validity (Valid): it reports the ratio of valid predicted SMILES representations that can be successfully parsed into a molecule. It is calculated among all the generated outputs that contain extractable answer part. If an output refuses to answer a question, it would not be counted for calculating the validity. Except for this metric, all the other metrics for SMILES are calculated based on the valid samples.
Additional metrics used in Appendix B are briefly introduced as follows:
@@
Appendix B Overall Experimental Results
B.1 Main Results
In this section, we show the overall experimental results on additional metrics. Results for name conversion tasks are presented in Tables 4,5,6, and 7. Results for molecule description tasks are presented in Tables 8 and 9. Results for chemical reaction tasks can be found in Tables 10 and 11. Results for the property prediction tasks can be found in Tables 13 and 12.
B.1.1 Name Conversion
For the NC-I2F task (Table 4), LlaSMol is the best performing LLM, tying with many other methods on validity. All LlaSMol models outperform all other methods except the SoTA task-specific method. This showcases the benefit of fine-tuning on SMolInstruct. The SoTA task-specific method (STOUT+RDKit) outperforms LlaSMol, achieving an exact match 97.9% of the time. The best performing LLM baseline for top-1 exact match is Galactica, and GPT-4 (1-shot) is the best performing LLM baseline for top-3. Neither of these methods come close to achieving the same performance as the LlaSMol series, except on Validity.
@@ B.1.1 Name Conversion
For the NC-I2S task (Table 5), LlaSMol is the best performing LLM for all metrics, tying with LlaSMol on validity. All LlaSMol models outperform all other methods except the SoTA task-specific method. This showcases the benefit of fine-tuning on SMolInstruct. The SoTA task-specific method (STOUT) only barely outperforms LlaSMol, achieving an exact match 73.5% of the time and even better FTS and validity scores. The best performing LLM baseline is Galactica, but its performance is not remotely close to LlaSMol for any metric except validity.
@@
| Molinst (instruction-tuned) | |||||||
@@
For the NC-S2F task (Table 6), LlaSMol is the best performing LLM for all metrics, tying with many other methods on validity. All LlaSMol models outperform all other methods except the SoTA task-specific method, which directly computes the answer using an algorithm and is therefore 100% accurate. The best performing LLM baselines are GPT-4 and GPT-4 (0 shot), but their performance is not remotely close to LlaSMol for any metric except validity. LlaSMol’s ability to translate SMILES representations to molecular formulas demonstrates an understanding of SMILES and chemical formulas. While this is not a challenging task, evaluating it can illustrate an LLM’s understanding of representations.
@@
For the NC-S2I task (Table 7), LlaSMol is the best performing LLM for all metrics. Interestingly, the models not fine-tuned on SMolInstruct cannot exactly match the expected output even 1% of the time. The SoTA task-specific method achieves 56.5% accuracy, indicating that this task is not necessarily easy. LlaSMol’s ability to translate SMILES representations to IUPAC names suggests a level of understanding of the functional groups in the IUPAC specification, as well as an understanding of SMILES representations.
@@
B.1.2 Molecule Description
For the MG task (Table 8), LlaSMol is the best performing LLM on all metrics except validity, where LlaSMol outperforms it by a small margin. Its top-1 exact match score is especially impressive, more than tripling the top-1 performance of the best baseline LLM (Molinst). This indicates that the right combination of foundation model and fine-tuning dataset can support a better comprehension of natural language about molecules.
@@ B.1.2 Molecule Description
For the MC task (Table 9), LlaSMol is the best performing LLM on all metrics. It is still outperformed by the SoTA task-specific model (MolT5), however, approaching close to it. Please note that these text-based metrics only measures the , and it does not necessary mean the correctness of the description on chemistry dimension. That is to say, these metrics are only a reference for how the description looks like the gold one, instead of a precise measure of correctness. Limited by the current resources, we cannot obtain the correctness measures, and will leave this to our future work.
| Model | BLEU-2 | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-L | METEOR |
| MolT5 | ||||||
| GPT-4 (0-shot) | ||||||
| GPT-4 (1-shot) | ||||||
| GPT-4 (3-shot) | ||||||
| GPT-4 (5-shot) | ||||||
| Llama 2 | ||||||
| Code Llama | ||||||
| Mistral | ||||||
| Galactica | ||||||
| Molinst (instruction-tuned) | ||||||
| LlaSMol | ||||||
| LlaSMol | ||||||
| LlaSMol | ||||||
| LlaSMol |
B.1.3 Chemical Reaction
@@
B.1.4 Property Prediction
| Model | ESOL | Lipo |
| Uni-Mol | ||
| GPT-4 (0-shot) | ||
| GPT-4 (1-shot) | ||
| GPT-4 (3-shot) | ||
| GPT-4 (5-shot) | ||
| Galactica | ||
| Llama 2 | ||
| Code Llama | ||
| Mistral | ||
| Molinst (instruction-tuned) | ||
| LlaSMol | ||
| LlaSMol | ||
| LlaSMol | ||
| LlaSMol |
For the two regression tasks in molecular property prediction (Table 12), LlaSMol achieves the highest performance among LLMs in both tasks. Notably, LlaSMol outperforms the best LLMs GPT-4 with a large margin, while all the other LlaSMol models also outperform it on Lipo task. This indicates the power of fine-tuning on SMolInstruct for understanding properties of molecules. However, the task-specific method (Uni-Mol) still outperforms all of the LLMs.
| Model | PP-BBBP | PP-ClinTox | PP-HIV | PP-SIDER | ||||||||
| F1 | MCC | Acc | F1 | MCC | Acc | F1 | MCC | Acc | F1 | MCC | Acc | |
| Uni-Mol | ||||||||||||
| GPT-4 (0-shot) | ||||||||||||
| GPT-4 (1-shot) | ||||||||||||
| GPT-4 (3-shot) | ||||||||||||
| GPT-4 (5-shot) | ||||||||||||
| Galactica | ||||||||||||
| Llama 2 | ||||||||||||
| Code Llama | ||||||||||||
| Mistral | ||||||||||||
| Molinst (instruction-tuned) | ||||||||||||
| LlaSMol | ||||||||||||
| LlaSMol | ||||||||||||
| LlaSMol | ||||||||||||
| LlaSMol | ||||||||||||
For the four classification tasks in molecular property prediction (Table 13), LlaSMol achieves the highest accuracy values among LLMs on all the tasks. Particularly, on PP-SIDER task, LlaSMol outperforms all the LLMs and the task-specific model Uni-Mol on all the three metrics, which highlights the potential of LLMs in understanding molecules and predicting their properties. However, LlaSMol achieves very low F1 values on PP-ClinTox and PP-HIV. This suggests that LlaSMol can struggle with data imbalance issue in these two tasks, and achieve high accuracy values by predicting most samples as negative. Similarly, most other LLMs also achieve either poor accuracy values (e.g., 6.3% and 4.4% for Molinst on PP-ClinTox and PP-HIV) or poor F1 values (e.g., 0.0% and 0.0% for Galactica). Therefore, there is still room for further improvement of the robustness of LLMs when dealing with imbalanced datasets.
B.2 Influence of LoRA Modules and Trainable Parameters
Figure 4 presents the model performance on all the 14 tasks under different LoRA settings and base models, as described in Section 5.4. As shown in Figure 4, one of the key observations is that incorporating more LoRA modules during training can consistently enhance the performance of LlaSMol on most tasks. In addition, by comparing LlaSMol and LlaSMol Large, another key observation we can tell is that LlaSMol Large with larger base model consistently outperforms all the LlaSMol models under different LoRA settings on almost all the tasks, except four classification tasks for molecular property prediction. These observations are consistent with those described in Section 5.4.
Appendix C More Statistics of SMolInstruct
To know more about the proposed SMolInstruct dataset, we do a statistics on the molecules (represented in SMILES). Altogether, it contains 1.6M distinct molecules, and several important statistical values are shown in results are shown Figure 5. Specifically, Bertz complexity is a topological index that measures the complexity of molecules based on the number and types of bonds and atoms. Atom count shows the number of atoms in a molecule, it represents the size of a molecule. Molecular weight is a measure of the sum of the atomic weights of the atoms in a molecule. And ring count shows the number of rings in the molecular structures. As we can see, the values varies much, showing a extensive coverage in terms of complexity, size, and structure. Notably, when compared to Mol-Instructions (Fang et al., 2023), molecules in SMolInstruct show a larger complexity and diversity, which indicates that tasks of SMolInstruct can be more comprehensive and complicated than those of Mol-Instructions. The scale, diversity, and careful construction of SMolInstruct makes it well-suited for learning chemistry LLMs.

### v4

@@ Appendix D Details of Experimental Setup
In this section, we introduce the details of our experimental setups, including the training and inference details of our LlaSMol models and the compared models. We also give detailed explanations of the metrics used in Section 4.3, as well extra metrics that we will use in Appendix E.
D.1 LlaSMol Models
@@
The base models used for developing LlaSMol are Galactica666https://huggingface.co/facebook/galactica-6.7b (Taylor et al., 2022), Llama 2777https://huggingface.co/meta-llama/Llama-2-7b-hf (Touvron et al., 2023b), Code Llama888https://huggingface.co/codellama/CodeLlama-7b-hf (Roziere et al., 2023) and Mistral999https://huggingface.co/mistralai/Mistral-7B-v0.1 (Jiang et al., 2023). We conduct instruction tuning on our SMolInstruct, and the resulting models are called named as LlaSMol, LlaSMol, LlaSMol, and LlaSMol, respectively. Expect for being based on different base models, their training and evaluation configurations are identical, as described as follows.
@@
During inference, we adopt beam search as the generation strategy for simplicity. Due to the need of evaluations on the top- predicted answers (as in Appendix E, where varies for different tasks, we generate different numbers of sequences for different tasks by setting the num_return_sequences argument in the Huggingface Transformers library (Wolf et al., 2020). Specifically, it is set to 5 for NC-I2S, NC-S2I, FS, and MG; 3 for NC-I2F and NC-S2F; 1 for all the PP tasks; and 10 for RS. The beam size is set to num_return_sequences + 3 for all the tasks. The maximum number of new generated tokens is set to 1024.
D.2 Compared LLMs
We introduce each of the compared LLMs in details, including their training (if applicable) and inference process.
D.2.1 GPT-4
@@
We conduct both -shot evaluations, where is the number of provided ICL examples. For -shot evaluation, the ICL part in the template is removed from the queries. In -shot evaluation, for each sample,the ICL examples are randomly selected from the training set. The results of these settings are shown in Appendix E, which reals that these settings’ performance is not consistent across all the tasks. Since -shot shows the best performance on most tasks, we report its results in Section 4.3.
In the evaluations, we use the default generation strategy set in the API. To generate the same number of results for each sample (as described in Section D.1), we set the argument n in the API, which controls the number of output sequences.
@@
D.2.2 Claude 3 Opus
Claude 3 Opus (Anthropic, 2024) is a newly proposed SoTA LLM to date. Similarly to GPT-4, we evaluate Claude 3 Opus on 500 samples from SMolInstruct test set via Anthropic’s API, and the generation strategy is the default one. The used prompt format is identical to the one used for GPT-4 (Section D.2.1. For each sample, we generate one response. Since Claude 3 Opus can always follow the formatted instructions, we do not bother to extract the answers from its outputs, but directly use its outputs as the predicted answers.
D.2.3 Galactica
@@
Galactica (Taylor et al., 2022) is a LLM without instruction tuning. To evaluate it on SMolInstruct, we follow the instructions in the paper (Taylor et al., 2022) and the repository101010https://github.com/paperswithcode/galai to create the queries for each task. We use zero-shot setting, as its official instruction does not suggest using few-shot setting. The generation configuration is set identical to that of our LlaSMol models (Section D.1).
@@
Galatica’s outputs may contain extra text other than the expected answers. Therefore, with heuristic rules and regular expression matching, we implement a program to extract the answers from the outputs of the models. Since the extraction cannot possibly cover all the possible output formats, some answers might not be correctly extracted, which might lead to validities lower than the actual value.
D.2.4 Llama 2, Code Llama, and Mistral
@@
For our base models (Llama 2, Code Llama, and mistral), since they are not trained on SMolInstruct and have not seen the diverse queries in the dataset, we use the same query templates as those used for GPT-4 (Section D.2.1). We use the one-shot setting for them, as it would improve models’ abiltity to follow the instructions and generate answers in a more formated way. In addition, the generation configuration (including beam size, output sequence numbers, etc) is set identical to that of our LlaSMol models (Section D.1).
@@
Although we try our best to make the output format as clear as possible in the queries, these three models still cannot follow the instructions and their outputs are in various formats. By heuristic rules and regular expression matching, we implement a program to extract the answers from the outputs of each of the models. Since the extraction cannot possibly cover all the possible output formats, some answers might not be correctly extracted, which might lead to validities lower than the actual value.
D.2.5 Molinst
@@
Molinst is a Llama 2 model fine-tuned on Mol-Instructions by Fang et al. (2023). On the shared tasks between Mol-Instructions and SMolInstruct (including MC, MG, FS, and RS), we directly use the query templates from Mol-Instructions to achieve better results. On other tasks, we create one query template for each task following the style of Mol-Instructions. We use zero-shot on its evaluation, as Mol-Instructions does not contain any few-shot use cases. We directly use Molinst’s checkpoint111111https://huggingface.co/zjunlp/llama2-molinst-molecule-7b. and evaluate it on SMolInstruct, which is a fair comparison, because SMolInstruct’s test set does not contain any training examples from the training set of Mol-Instructions.
@@
The outputs of Molinst may also contain extra text other than the expected answers, especially on its unseen tasks. Thus, we also implement a program to extract the answers. Since the extraction cannot possibly cover all the possible output formats, some answers might not be correctly extracted, which might lead to their validities lower than the actual value.
D.2.6 ChemLLM
ChemLLM (Zhang et al., 2024) is an LLM simultaneously proposed with ours. We apply their released checkpoint121212https://huggingface.co/AI4Chem/ChemLLM-7B-Chat and evaluate it on SMolInstruct. The used prompt format is identical to the one used for GPT-4 (Section D.2.1. For each sample, we generate one response.
The outputs of ChemLLM may contain extra text other than the expected answers. Thus, we implement a program to extract the answers from its outputs. Since the extraction cannot possibly cover all the possible output formats, some answers might not be correctly extracted, which might lead to their validities lower than the actual value.
D.3 Task-Specific, Non-LLM Based SoTA Models
D.3.1 STOUT for NC-I2S and NC-S2I
@@
STOUT is a encoder-decoder model trained on SMILES and IUPAC name paired data, and it is capable of conducting the NC-2iS and NC-S2I tasks. Due to the lack of training code, we cannot re-train it on our dataset, and directly use their released model checkpoint131313https://github.com/Kohulan/Smiles-TO-iUpac-Translator. Since it may have encounter some test samples of SMolInstruct during training, the evaluation results in Table 1 may be higher than its real performance.
D.3.2 RDKit for NC-S2F
@@ D.4 Evaluation Metrics
Fingerprint Tanimoto Similarity (FTS). It is an important metric type commonly used in cheminformatics. It measures the structural similarity between molecules. The one we report in Section 4.3 is one of this type, called Morgan FTS, which leverages Morgan method to calculate the fingerprint (Morgan, 1965).
@@
Validity (Valid): it reports the ratio of valid predicted SMILES representations that can be successfully parsed into a molecule. It is calculated among all the generated outputs that contain extractable answer part. If an output refuses to answer a question, it would not be counted for calculating the validity.
Additional metrics used in Appendix E are briefly introduced as follows:
@@
Appendix E Detailed Experimental Results
In this section, we present more comprehensive experimental results and analysis. We will provide detailed performance of all models across a wider range of metrics and offer more discussion of the findings. We will discuss about each type of tasks in the rest of this section.
E.1 Name Conversion Tasks
The results on the four NC tasks are presented in Table 7, Table 8, Table 9, and Table 10. On these tasks, Although open source LLMs including Llama 2, Code Llama, and Mistral can sometimes achieve fairly good validity, showing that they know at least basic knowledge of the molecular representations, they correctly predict none of the samples. It indicates that their lack of knowledge to do reasoning for the conversions. Trained on scientific corpus (including chemistry materials), Galactica can achieve some correct predictions, but the performance is still low. After fine-tuning on these open source LLMs, LlaSMol models are substantially better than other LLMs, showing the benefits of our dataset and fine-tuning. Compared to the SoTA models, LlaSMol models still underperform, but approach them on NC-I2F, NC-I2S, and NC-S2F. However, on NC-S2I, LlaSMol is still far from SoTA. It seems that NC-S2I is the hardest task among the four NC tasks, as the SoTA task-specific method, STOUT, only achieves 56.5% accuracy. Ability on this task suggests a level of understanding of the functional groups in the IUPAC specification, as well as an understanding of SMILES representations. Improving models’ understanding and generation for IUPAC names might be necessary.
@@ E.1 Name Conversion Tasks
| Mistral | |||||||
| Molinst | |||||||
| ChemLLM | - | - | |||||
@@
E.2 Property Prediction
| Model | ESOL | Lipo |
| Uni-Mol | ||
| GPT-4 (0-shot) | ||
| GPT-4 (1-shot) | ||
| GPT-4 (3-shot) | ||
| GPT-4 (5-shot) | ||
| Claude 3 | ||
| Galactica | ||
| Llama 2 | ||
| Code Llama | ||
| Mistral | ||
| Molinst | ||
| ChemLLM | ||
| LlaSMol | ||
| LlaSMol | ||
| LlaSMol | ||
| LlaSMol |
The results on the two regression tasks (PP-ESOL and PP-Lipo) are presented in Table 11. LlaSMol outperforms other LLMs including GPT-4 by a large margin, showing the effectiveness of the fine-tuning. However, it is slightly worse than Claude 3 Opus on PP-ESOL, possibly because of its much larger scale of parameters and training data. The task-specific method (Uni-Mol) still outperforms all of the LLMs. This might be attributed to the fact that Uni-Mol (Zhou et al., 2023) possesses 3D structure knowledge, which may be useful in predicting molecular properties.
| Model | BBBP | ClinTox | HIV | SIDER |
| Uni-Mol | ||||
| GPT-4 (0-shot) | ||||
| GPT-4 (1-shot) | ||||
| GPT-4 (3-shot) | ||||
| GPT-4 (5-shot) | ||||
| Claude 3 Opus | ||||
| Galactica | ||||
| Llama 2 | ||||
| Code Llama | ||||
| Mistral | ||||
| Molinst | ||||
| ChemLLM | ||||
| LlaSMol | ||||
| LlaSMol | ||||
| LlaSMol | ||||
| LlaSMol |
The results for the four classification tasks (PP-BBBP, PP-Clintox, PP-HIV, and PP-SIDER) presented in (Table 12) show similar information. Particularly, on PP-SIDER task, LlaSMol outperforms all the LLMs and the task-specific model Uni-Mol, which highlights the potential of LLMs in understanding molecules and predicting their properties. The reason why LlaSMol models get the same results on PP-HIV is that the dataset is imbalanced, and the models all predict most samples as negative samples.
E.3 Molecule Description
For the MC task (Table 13), LlaSMol is the best performing LLM on all metrics. It is still outperformed by the SoTA task-specific model (MolT5), however, approaching close to it. Please note that these text-based metrics only measures the similarity to the gold descriptions, and it does not necessary mean the correctness of the description on chemistry dimension. Limited by the current resources, we cannot obtain the correctness measures, and will leave this to our future work.
| Model | BLEU-2 | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-L | METEOR |
| MolT5 | ||||||
| GPT-4 (0-shot) | ||||||
| GPT-4 (1-shot) | ||||||
| GPT-4 (3-shot) | ||||||
| GPT-4 (5-shot) | ||||||
| Claude 3 Opus | ||||||
| Galactica | ||||||
| Llama 2 | ||||||
| Code Llama | ||||||
| Mistral | ||||||
| Molinst | ||||||
| ChemLLM | ||||||
| LlaSMol | ||||||
| LlaSMol | ||||||
| LlaSMol | ||||||
| LlaSMol |
For the MG task (Table 14), LlaSMol is the best performing LLM on all metrics. Since this task aims to generate an arbitrary molecule that fits the given description, it is not required to exactly match the gold one, and the FTS metrics may be better to reflect models’ ability. However, similar to MC, the metrics only measure the similarities between predicted and gold ones, and not necessarily means the correctness of the generation. We will leave this to our future work.
@@ E.3 Molecule Description
E.4 Chemical Reaction
@@
E.5 Other Common Findings
Besides for the aforementioned findings, we can observe some common findings across different tasks. Firstly, we can see that GPT-4 does not show consistent pattern on different in-context learning settings, with 0-shot achieving the best performance on most tasks. It may indicate that GPT-4 does not have sufficient knowledge about chemistry and cannot effectively learn how to do these tasks by imitating several samples. Injecting chemistry knowledge by training LLMs may be necessary. Secondly, although Claude 3 Opus still trails behind LlaSMol and the SoTA task-specific models, it surpasses GPT-4 on all the chemistry tasks. Claude 3 Opus demonstrates notably enhanced performance compared to GPT-4, showcasing its superior grasp of chemistry knowledge, highlighting its potential applications in the domain.
Appendix F More Analytical Experiments
F.1 Task Synergy
In order to investigate the synergy among different tasks, we first conduct an experiment to evaluate the performance of an LLM trained on single task or a specific type of task. We train several single-task models, each focusing on one task or a group of tasks from the same type141414These single-task models include LlaSMol-NC, LlaSMol-PP, LlaSMol-MC, LlaSMol-MG, LlaSMol-FS, and LlaSMol-RS, where the model names indicate the training task(s). For instance, LlaSMol-FS is trained only on the FS task, LlaSMol-NC is trained on all the NC tasks combined, and LlaSMol-PP is trained on all the PP tasks combined. The training setups are identical to those of LlaSMol.. We then compare the performance of these single-task models with LlaSMol, which is trained on all 14 tasks, to assess the potential benefits of multi-task training.
| Task | Metric | Single-Task | Multi-Task | Improv. |
| NC-I2F | EM (%) | 86.8 | 87.9 | 1.1 |
| NC-I2S | EM (%) | 67.6 | 70.1 | 2.4 |
| NC-S2F | EM (%) | 93.2 | 93.2 | 0.0 |
| NC-S2I | EM (%) | 27.4 | 29.0 | 1.5 |
| PP-ESOL | RMSE | 20.616 | 1.150 | 19.466 |
| PP-Lipo | RMSE | 1.241 | 1.010 | 0.231 |
| PP-BBBP | Acc (%) | 68.5 | 74.6 | 6.1 |
| PP-Clintox | Acc (%) | 79.9 | 93.1 | 13.2 |
| PP-HIV | Acc (%) | 96.7 | 96.7 | 0.0 |
| PP-SIDER | Acc (%) | 64.3 | 70.7 | 6.4 |
| MC | METEOR | 0.299 | 0.452 | 0.153 |
| MG | FTS (%) | 33.1 | 61.7 | 28.6 |
| FS | EM (%) | 62.6 | 63.3 | 0.7 |
| RS | EM (%) | 31.5 | 32.9 | 1.4 |
The results are presented in Table 17. The multi-task trained model, LlaSMol, outperforms the single-task models on the majority of tasks. The improvements are particularly substantial on PP-ESOL, PP-BBBP, PP-Clintox, MC, and MG. This finding suggests the presence of shared knowledge across these tasks. For instance, the knowledge of understanding SMILES may be common to most tasks, and training models on multiple tasks can enhance this knowledge, thus achieving better performance.
To further investigate the relationships among different tasks, we conduct an additional experiment by removing certain tasks from the training data. Specifically, we train a set of models, each with one task or a group of tasks from the same type removed from its training data151515These models are named as “w/o Task”, where Task represents the task names., and their base models and training setups are identical to those of LlaSMol. We then compare their performance with LlaSMol, which is trained without removing any tasks.
| Model | NC-I2F | NC-I2S | NC-S2F | NC-S2I | PP-ESOL | PP-Lipo | PP-BBBP | PP-Clintox | PP-HIV | PP-SIDER | MC | MG | FS | RS |
| EM (%) | EM (%) | EM (%) | EM (%) | RMSE | RMSE | Acc (%) | Acc (%) | Acc (%) | Acc (%) | METEOR | FTS (%) | EM (%) | EM (%) | |
| w/o NC | - | - | - | - | 1.520 | 1.090 | 76.1 | 93.1 | 96.8 | 70.6 | 0.436 | 54.9 | 63.2 | 33.5 |
| w/o PP | 87.9 | 70.7 | 93.5 | 28.7 | - | - | - | - | - | - | 0.447 | 62.3 | 64.2 | 33.1 |
| w/o MC | 87.6 | 71.0 | 93.5 | 27.8 | 1.133 | 1.057 | 74.1 | 93.1 | 96.8 | 70.9 | - | 64.1 | 63.3 | 33.4 |
| w/o MG | 87.8 | 69.6 | 93.4 | 27.8 | 1.231 | 0.982 | 77.2 | 93.1 | 96.8 | 70.9 | 0.445 | - | 63.4 | 34.0 |
| w/o FS | 87.9 | 70.4 | 93.8 | 29.5 | 1.278 | 1.288 | 70.6 | 93.1 | 96.8 | 70.8 | 0.452 | 63.2 | - | 33.1 |
| w/o RS | 88.0 | 71.1 | 93.7 | 29.7 | 1.203 | 1.048 | 72.1 | 93.1 | 96.8 | 70.6 | 0.450 | 62.6 | 61.9 | - |
| LlaSMol | 87.9 | 70.1 | 93.2 | 29.0 | 1.150 | 1.010 | 74.6 | 93.1 | 96.7 | 70.7 | 0.452 | 61.7 | 63.3 | 32.9 |
Interestingly, results presented in Table 18 generally show no consistent pattern among different tasks. Specifically, removing certain task(s) can lead to improvements on some tasks, while leading to decrements on others. And in most cases, the change is not substantial, suggesting that each task does not heavily rely on any other tasks, but rather on the data of the task itself. However, there are two exceptions – for the “w/o NC” model, the performance of PP-ESOL and MG greatly drops. On PP-ESOL, we hypothesize that the knowledge of understanding chemical structures learned on NC might be useful in predicting the solubility of a molecule. For MG, since there are many IUPAC names in the input molecular descriptions, the performance drop may be attributed to the fact that the ability to convert IUPAC names to SMILES, learned on NC, can help the model directly obtain the related SMILES representations, thus leading to better performance.
F.2 Influence of LoRA Modules and Trainable Parameters
In this section, we investigate the influence of using different LoRA modules, or different sizes of trainable parameters. We take LlaSMol as the basic setting and refer to it as LlaSMol in this section for simplicity. All the compared models are listed as follows, with trainable parameter sizes and ratios labeled in brackets:
LlaSMol Lite (6.8M, 0.09%): LoRA is applied on q_proj and v_proj of the attention modules.
LlaSMol Attn (13.6M, 0.19%): LoRA is applied on all the attention projection matrices (including q_proj, k_proj, v_proj, o_proj).
LlaSMol FFN (28.3M, 0.39%): LoRA is applied on all the FFN projection matrices (including gate_proj, down_proj, up_proj).
LlaSMol (41.9M, 0.58%): The basic setting. LoRA is applied on all the attention and FFN projection matrices.
LlaSMol Plus (173.0M, 2.33%): LoRA is applied on all the attention and FFN projection matrices, and lm_head is set trainable.
Figure C presents the model performance on all the 14 tasks. We can observe that for most tasks, progressing from LlaSMol Lite to LlaSMol Attn, LlaSMol FFN, and finally LlaSMol, the incorporation of more LoRA modules (and thus more trainable parameters) leads to a substantial performance enhancement. Further adding lm_head as trainable parameters in LlaSMol Plus can further slightly improves the performance. This indicates that refining the selection of LoRA modules and incorporating more trainable parameters is very important. Moreover, LLMs exhibit considerable potential to surpass the performance of previous task-specific models on various chemistry tasks, provided that more trainable parameters are allowed or even full fine-tuning is employed.
However, for certain property prediction (PP) tasks, the incorporation of additional LoRA modules does not demonstrate consistent improvements. This inconsistency may be attributed to the relatively small number of training and test samples, which can result in reduced model robustness when faced with diverse inputs, as well as the inherent randomness associated with limited data.